A string is stretched and fastened to two points x = 0 and x = l apart. Motion is started by



displacing the string into the form y = k(lx − x2



) from which it is released at time t = 0. Find



the displacement of any point on the string at a distance of x from one end at time t.



Hint: From this problem, we have the following boundary conditions:



y(0,t) = 0 for all t > 0



y(l,t) = 0 for all t > 0



∂y



∂t



(x, 0) = 0 (initial velocity is zero)



y(x, 0) = k(lx − x2



)

1
Expert's answer
2021-11-30T12:49:00-0500

"y=k(lx-x^2)"

Boundary conditions:

1. "y(0,t)=0" For all t>0

2. "y(1,t)=0" For all t>0

3. "\frac{dy}{dt}(y,0)=0"

4. "y(x,0)=k(lx-x^2)"

The displacement y(x,t) is given by the equation

"\frac{d^2y}{dt^2}=a^2\frac{d^2y}{dx^2}" ..........(i)

The suitable solution of (i) is given by

"y(x,t)=(Acos\ lx+Bsin\ lx)(Ccos\ lat+Dsin\ lat)"......(ii)

Using the first and second boundary conditions in (ii) we get,

A=0 and "\lambda=\frac{nπ}{l}"

"\therefore y(x,t)=Bsin\frac{nπx}{l}(Ccos \frac{nπat}{l}+Dsin\frac{nπat}{l}.........(iii)"

Now "\frac{dy}{dt}=Bsin\frac{nπx}{l}(-Csin\frac{nπat}{l}\cdot\frac{nπa}{l}+Dcos\frac{nπat}{l}\cdot\frac{nπa}{l})"

Using third boundary condition,

0="Bsin\frac{nπx}{l}D\frac{nπa}{l}"

Here, B can not be 0, hence,D=0

Thus equation (iii) becomes;

"y(x,t)=Bsin\frac{nπx}{l}\cdot Ccos\frac{nπat}{l}"

"=B_1sin\frac{nπx}{l}\cdot cos\frac{nπat}{l}" ,where B1=Bc

The most general solution is;

"y(x,t)=\displaystyle\sum_{n=1}^{\infty}Bsin\frac{nπx}{l} cos\frac{nπat}{l}" ....(iv)

Using the forth boundary condition;

"k(lx-x^2)=\displaystyle\sum_{n=1}^{\infty}Bsin\frac{nπx}{l}" .....(v)

The RHS of (v) is the half range Fourier sine series of the LHS function

"\therefore B_n=\frac{2}{l}\int_0^lf(x)\cdot sin\frac{nπx}{l}dx"

"=\frac{2k}{l}\int_0^l(lx-x^2)d(\frac{-cos\frac{nπx}{l}}{\frac{nπ}{l}})"

"=\frac{2k}{l}[(lx-x^2)d(\frac{-cos\frac{nπx}{l}}{\frac{nπ}{l}})-(l-2x)(\frac{-sin\frac{nπx}{l}}{\frac{n^2π^2}{l^2}})+(-2)(\frac{cos\frac{nπx}{l}}{\frac{n^3π^3}{l^3}}]_0^l"

"=\frac{2k}{l}(\frac{-2cosnπ}{\frac{n^3π^3}{l^3}})+(\frac{2}{\frac{n^3π^3}{l^3}})"

"=\frac{2k}{l}\cdot\frac{2l^3}{n^3π^3}(1-cosnπ)"

"i.e, B_n=\frac{4kl^2}{n^3π^3}(1-(-1)^n)\ or\ B_n= \begin{cases}
 \frac{8kl^2}{n^3π^3} &\text{if } n\ is\ odd \\
 0 &\text{if } n\ is\ even
\end{cases}"

"\therefore y(x,t)=\displaystyle\sum_{n=odd}^{\infty}\frac{8kl^2}{n^3π^3}cos\frac{nπat}{l}sin\frac{nπx}{l}"

Or "y(x,t)=\frac{8k}{π^3}\displaystyle\sum_{n=1}^{\infty}\frac{1}{(2n-1)^3}cos\frac{(2n-1)πat}{l}sin\frac{(2n-1)πx}{l}"